# Bounding the degree of generic sharp transitivity

### v1

@@ Abstract.
We show that a generically sharply $t$-transitive permutation group of finite Morley rank on a set of rank $r$ satisfies $t\leq r+2$ provided the pointwise stabilizer of a generic $(t-1)$-tuple is an $L$-group, which holds, for example, when the stabilizer is solvable or when $r\leq 5$. This makes progress on the Borovik-Cherlin conjecture that every generically $(r+2)$-transitive permutation group of finite Morley rank on a set of rank $r$ is of the form $\operatorname{PGL}_{r+1}(F)$ acting naturally on $\mathbb{P}^{r}(F)$.
@@ 1. Introduction
The study of generically $t$-transitive actions was first taken up by Popov in the context of algebraic groups [Pop07] and later generalized to the model-theoretic setting of groups of finite Morley rank by Borovik and Cherlin [BC08]. While ordinary $t$-transitivity requires that a group $G$ act transitively on $X^{t}$ off of the diagonal, generic $t$-transitivity only requires $G$ act transitively on $X^{t}$ off of some set of smaller dimension.
@@ 1.1. The problem
It is a classical result that ordinary $t$-transitivity does not exist for actions on an infinite set once $t\geq 4$ [Tit52, Hal54], but as we have already observed, generic $t$-transitivity exists very naturally for all $t$. However, one may still look to bound the degree of generic transitivity relative to a fixed $G$ or for actions on sets of a fixed rank.
@@ 1.2. The theorem
At present, little has been done around Problem 1, but its solution for modules provides a sliver of hope. Earlier work of the authors addressed aspects of Problem 3 in arbitrary rank [AW19]; here, we study Problem 2, which previously was only known when $r\leq 2$ as consequence of [AW18] (see also the early work of Gropp [Gro92]).
Our main result is rank free but comes with a condition on the stabilizer of a generic $(t-1)$-tuple from $X$. In the target case when $(G,X)\cong(\operatorname{PGL}_{r+1}(F),\mathbb{P}^{r}(F))$, this stabilizer is precisely a maximal torus of $G$, and for our result, we “simply” ask that the stablizer does not involve a certain (conjecturally nonexistent) class of simple groups. In particular, the condition will automatically be met if this stabilizer is known to be solvable. The precise framework we work in is that of $L$-groups.
@@
  • $H$ has even type if contains an infinite elementary abelian $2$-group but no Prüfer $2$-group;
  • $H$ has odd type if contains a Prüfer $2$-group but no infinite elementary abelian $2$-group;
  • $H$ has mixed type if contains an infinite elementary abelian $2$-group and a Prüfer $2$-group.
@@
Assume $(G,X)$ is a generically sharply $t$-transitive permutation group of finite Morley rank with $\operatorname{rk}X=r$. Let $H$ be the pointwise stabilizer of a generic $(t-1)$-tuple, and suppose that $H$ is an $L$-group.
If $t\geq r+2$, then $t=r+2$, and $H$ is solvable.
We stress that the $L$-hypothesis is not placed on $G$ but rather only on $H$, which conjecturally should be a torus. And as mentioned above, it is [AW24] that provides traction for the $L$-hypothesis; in particular, [AW24, Corollary] tells us at the outset that either $H$ is solvable or contains involutions.
@@ 1.3. The strategy
Assume we are in the setting of the Theorem with $t\geq r+2$, $(1,\ldots,t)$ a generic $t$-tuple in $X^{t}$, and $G_{[t-1]}$ the pointwise stabilizer of $11,\ldots,t-$ 1. Let $\Sigma$ be the setwise stabilizer of $1,\ldots,t$ and $\Sigma_{t}\leq\Sigma$ the stabilizer of $t$ in $\Sigma$. Generic sharp $t$-transitivity readily implies $\Sigma\cong\operatorname{Sym}(t)$, $\operatorname{rk}G_{[t-1]}=r$, and $\Sigma_{t}$ acts faithfully on $G_{[t-1]}$, see [AW18, Section 4.4].
@@
The main point then is that carrying a faithful action of a large symmetric group is quite restrictive. In Section 2.1 we study actions of $\operatorname{Alt}(n)$ and $\operatorname{Sym}(n)$ on connected nonsolvable $L$-groups of rank $r$ and show in Proposition 1 that typically $n\leq r$. In the setting of the Theorem, this yields that $G_{[t-1]}$ is either solvable, or we are in the exceptional case of $\Sigma_{t}\cong\operatorname{Sym}(5)$ and $G_{[t-1]}\cong\operatorname{PGL}_{2}(F)$ for $F$ a field of characteristic $5$ (which is an expected complication due the isomorphism $\operatorname{Sym}(5)\cong\operatorname{PGL}_{2}(\mathbb{F}_{5})$).
This leaves two quite different tasks: (1) prove that solvability of $G_{[t-1]}$ implies $t\leq r+2$, and (2) prove that $G_{[t-1]}\cong\operatorname{PGL}_{2}(F)$ does not survive. Both tasks require us to work with a larger chunk of $G$ than just $G_{[t-1]}$, but it turns out that $G_{[t-2]}$ (i.e. the pointwise stabilizer of $21,\ldots,t-$ 2) suffices.
@@
In the case when $G_{[t-1]}$ is solvable, the condition $t>r+2$ forces $G_{[t-1]}$ to be an elementary abelian $p$-group (using [AW24, Lemma 2.7] and [CDW23]); this configuration is then eliminated by Proposition 2, which looks generally at generically $2$-transitive actions with abelian point stabilizers. The case of $G_{[t-1]}\cong\operatorname{PGL}_{2}(F)$ forces $\operatorname{rk}G_{[t-2]}=6$, and to analyze this, we undertake a rather general study of simple groups of rank $6$ in Section 2.3—culminating with Proposition 3 showing that such groups are so-called $N_{\circ}^{\circ}$-groups in the sense of [DJ16]—which we then apply to $G_{[t-2]}$ to obtain a contradiction.
@@
We emphasize that the key ingredients of our proof—Propositions 1, 2, and 3—live beyond the specific context of the Theorem and should be interesting in their own right. We also note that the $L$-hypothesis in the Theorem is only used to apply Proposition 1. We would very much like to see the hypothesis removed from Proposition 1 (hence from the Theorem), but this seems considerably out of reach at present.
@@ 2.1. Actions of $\operatorname{Alt}(n)$ on nonsolvable groups
Leveraging the algebraic theory, we solve the conjecture for actions on $L$-groups; this is Proposition 1. Our proof consists to a large degree of simply collecting and organizing known results. As indicated in Corollary 2.5, the bound of $\operatorname{rk}H$ on $n$ is typically rather weak and more on the order of $\sqrt{\operatorname{rk}H}$.
@@
If $n\geq 9$, then $n\leq\operatorname{rk}H$. If the acting group is $\operatorname{Sym}(n)$, then $n\geq 5$ also implies $n\leq\operatorname{rk}H$ unless $n=5$ and $H=\operatorname{PGL}_{2}(F)$ with $\operatorname{char}F=5$.
@@
By [ABC08] (generalizing [Hum75, Theorem 27.4]), the acting group is contained in $\operatorname{Inn}(G)\cdot\Gamma$ for $\Gamma$ the group of graph automorphisms relative to a fixed maximal torus and Borel subgroup, so $\operatorname{Alt}(n)$ embeds into $G$ in the cases under consideration. By Corollary 2.5, it only remains to address when $5\leq n\leq 8$ and the acting group is $\operatorname{Sym}(n)$. Let $r=\dim G$. If $r\geq 8$, there is nothing to show, so we may assume $G$ is of type $A(1)$. In particular, $r=3$, and $G=\operatorname{PGL}_{2}(F)$ has no graph automorphisms. Thus, $\operatorname{Sym}(n)$ embeds into $\operatorname{PGL}_{2}(F)$. This is now quite classical. Via the adjoint representation of $\operatorname{PGL}_{2}(F)$, we have a faithful $3$-dimensional representation of $\operatorname{Sym}(n)$, so the only possibility is $n=5=\operatorname{char}F$ (see for example [Dic08]). ∎
@@
Since the the Morley rank of an algebraic group $G$ is always greater than or equal to its algebraic dimension, the bound $n\leq\dim G$ appearing in Corollary 2.6 can be replaced by $n\leq\operatorname{rk}G$.
@@
We will denote by $X$ the acting group, $n$ the degree of the permutation group $X$, and $r=\operatorname{rk}(H)$. Also let $X^{\prime}=[X,X]=\operatorname{Alt}(n)$.
We proceed by contradiction. Thus $r<n$, and we choose $H$ to be a counterexample of minimal rank $r$.
@@
If not, then $\operatorname{rk}(H/\sigma^{\circ}(H))<r$. Since $X$ acts on $H/\sigma^{\circ}(H)$, minimality of $r$ implies that this action is not faithful. As $X^{\prime}$ is simple and $[X^{\prime},H]$ is connected, we have $[X^{\prime},H]\leq\sigma^{\circ}(H)$. Now consider the action of $X$ on $\sigma^{\circ}(H)$. If $[X^{\prime},\sigma^{\circ}(H)]=1$, then $[X^{\prime},[X^{\prime},H]]=1$, and the three subgroups lemma implies that the group $[X^{\prime},X^{\prime}]=X^{\prime}$ centralizes $H$, against faithfulness of the action. Thus, $[X^{\prime},\sigma^{\circ}(H)]\neq 1$, so $X$ acts faithfully on $\sigma^{\circ}(H)$. Then [AW24, Lemma 2.7] implies $\operatorname{rk}(\sigma^{\circ}(H))\geq n-2$ in all cases except possibly when $X=\operatorname{Alt}(9)$. However, if $X=\operatorname{Alt}(9)$ and $\operatorname{rk}(\sigma^{\circ}(H))<n-2$, we can use that the minimal rank of a faithful connected $\operatorname{Alt}(9)$-module is $8$ (see the final remark of Section 3.3 in [CDW23]) to repeat the analysis of [AW24, Lemma 2.7] and eliminate this case. Thus $\operatorname{rk}(\sigma^{\circ}(H))\geq n-2$, meaning that $\operatorname{rk}(H/\sigma^{\circ}(H))\leq 2$. Since all connected groups of Morley rank at most $2$ are solvable, this is a contradiction, so $\sigma^{\circ}(H)=1$. ∎
@@
Notice that $H/\sigma(H)$ is also a counterexample of rank $r$, so the socle $\operatorname{soc}(H/\sigma(H))$ is as well. Thus we may assume $H=\operatorname{soc}(H/\sigma(H))$, which is semisimple by [BN94, Theorem 7.8]. ∎
@@
Since $H$ is an $L$-group, each $H_{i}$ is either a simple linear algebraic over an algebraically closed field or of degenerate type. By the inductive assumption, $n>r\geq k$ (in fact, $n>r\geq 3k$ since the $H_{i}$ are not solvable). Since $X^{\prime}$ is simple and permutes the set $\{1,\ldots,k\}$, the only possible action is the trivial one. Thus $X^{\prime}$ normalizes each $H_{i}$. However, $\operatorname{rk}H_{i}\leq r<n$, so $X^{\prime}$ centralizes each $H_{i}$ using Corollary 2.6 in the algebraic case and [AW24] in the degenerate case. This contradicts faithfulness. ∎
@@
$X=\operatorname{Sym}(n)$ with $5\leq n\leq 8$ yields a contradiction.
@@
By the contradictory assumption, $r\leq 7$. Since connected groups of Morley rank at most $2$ are solvable, $k$ is at most $2$. We first eliminate the case when $k=2$. If $k=2$ then one of the groups, say $H_{1}$ is of rank $3$ and the other is of rank $4$. As a result $X$ cannot permute these two groups and normalizes each one. Minimality of $r$ implies that the action of $X$ on both components has a nontrivial kernel, which much be $X^{\prime}$ by the subgroup structure of the symmetric group. It follows that the action of $X$ on $H$ is not faithful, a contradiction. Thus, $k=1$ and $H$ is simple. As in the previous claim, Corollary 2.6 and [AW24] provide the final contradiction. ∎

### v2

@@ Abstract.
We show that a generically sharply $t$-transitive permutation group of finite Morley rank on a set of rank $r$ satisfies $t\leq r+2$ provided the pointwise stabilizer of a generic $(t-1)$-tuple is an $L$-group, which holds, for example, when this stabilizer is solvable or when $r\leq 5$. This makes progress towards establishing the natural bound on $t$ implied by the Borovik-Cherlin conjecture that every generically $(r+2)$-transitive permutation group of finite Morley rank on a set of rank $r$ is of the form $\operatorname{PGL}_{r+1}(F)$ acting naturally on $\mathbb{P}^{r}(F)$.
@@ 1. Introduction
The study of generically $t$-transitive actions was popularized by Popov in the context of algebraic groups [Pop07] and then by Borovik and Cherlin in the more general setting of groups of finite Morley rank [BC08], although the first substantial results appeared nearly two decades earlier with work of Gropp [Gro92]. While ordinary $t$-transitivity requires that a group $G$ act transitively on $X^{t}$ off of the diagonal, generic $t$-transitivity only requires $G$ act transitively on $X^{t}$ off of some set of smaller dimension.
@@ 1.1. The problem
As we have already observed, generic $t$-transitivity exists very naturally for all $t$; however, one may still look to bound $t$ relative to a fixed $G$ or relative to actions on sets of a fixed rank.
@@ 1.2. The theorem
At present, little has been done around Problem 1, but its solution for modules provides a sliver of hope. Earlier work of the authors addressed aspects of Problem 3 in arbitrary rank [AW19]; here, we study Problem 2, which previously was only handled when $r\leq 2$ as consequence of [AW18] (see also the early work of Gropp [Gro92]).
Our main result is rank free but comes with a condition on the stabilizer of a generic $(t-1)$-tuple from $X$. In the target case when $(G,X)\cong(\operatorname{PGL}_{r+1}(F),\mathbb{P}^{r}(F))$, this stabilizer is precisely a maximal torus of $G$, and for our result, we “simply” ask that the stabilizer does not involve a certain (conjecturally nonexistent) class of simple groups. In particular, the condition will automatically be met if this stabilizer is known to be solvable. The precise framework we work in is that of $L$-groups.
@@
  • $H$ has even type if it contains an infinite elementary abelian $2$-group but no Prüfer $2$-group;
  • $H$ has odd type if it contains a Prüfer $2$-group but no infinite elementary abelian $2$-group;
  • $H$ has mixed type if it contains an infinite elementary abelian $2$-group and a Prüfer $2$-group.
@@
Assume $(G,X)$ is a generically sharply $t$-transitive permutation group of finite Morley rank with $\operatorname{rk}X=r$. Let $H$ be the pointwise stabilizer of a generic $(t-1)$-tuple. If $H$ is an $L$-group, then $t\leq r+2$.
We stress that the $L$-hypothesis is not placed on $G$ but rather only on $H$, which conjecturally should be a torus when $t\geq r+2$. As mentioned above, it is [AW24] that provides traction for the $L$-hypothesis; in particular, [AW24, Corollary] tells us at the outset that either $H$ is solvable or contains involutions.
Remark.
Our work also yields information when $t=r+2$. In this extremal case, Propositions 1 and 5 can be used to show that a generic $(t-1)$-tuple must necessarily be solvable except possibly when $r=4$. Analysis similar to what we undertake for $r=3$ in Proposition 5 (leveraging the significant Proposition 3) may establish solvability when $r=4$, but we do not take it up here.
@@ 1.3. The strategy
Let us outline our approach to the Theorem. Assume we are in the setting of the Theorem with $(1,\ldots,t)$ a generic $t$-tuple in $X^{t}$ and $G_{[t-1]}$ the pointwise stabilizer of $11,\ldots,t-$ 1. Let $\Sigma$ be the setwise stabilizer of $1,\ldots,t$ and $\Sigma_{t}\leq\Sigma$ the stabilizer of $t$ in $\Sigma$. Generic sharp $t$-transitivity readily implies $\Sigma\cong\operatorname{Sym}(t)$ and $G_{[t-1]}$ is connected of rank $r$. By [AW18, Lemma 4.27], $\Sigma_{t}$ acts faithfully on $G_{[t-1]}$. Towards a contradiction, we assume that $t\geq r+3$. The main point then is that $G_{[t-1]}$ carries a faithful action of the relatively large symmetric group $\operatorname{Sym}(r+2)$, and this is restrictive.
@@
In Section 2.1, we study actions of $\operatorname{Alt}(n)$ and $\operatorname{Sym}(n)$ on connected nonsolvable $L$-groups of rank $r$ and show in Proposition 1 that typically $n\leq r$. When $t\geq r+3$, this yields that either: (1) $G_{[t-1]}$ is solvable, or (2) we are in the exceptional case of $\Sigma_{t}\cong\operatorname{Sym}(5)$ and $G_{[t-1]}\cong\operatorname{PGL}_{2}(F)$ with $\operatorname{char}F=5$ (which is an expected complication as $\operatorname{Sym}(5)\cong\operatorname{PGL}_{2}(\mathbb{F}_{5})$). To push further, both cases require us to work with a larger chunk of $G$ than just $G_{[t-1]}$, but it turns out that $G_{[t-2]}$ (the pointwise stabilizer of $21,\ldots,t-$ 2) suffices.
@@
When $G_{[t-1]}$ is solvable, the condition $t\geq r+3$ forces $G_{[t-1]}$ to be an elementary abelian $p$-group. This configuration is studied at the level of $G_{[t-2]}$ and eliminated by Proposition 2, which looks at generically $2$-transitive actions with abelian point stabilizers. The case of $G_{[t-1]}\cong\operatorname{PGL}_{2}(F)$ corresponds to $\operatorname{rk}G_{[t-2]}=6$, and to analyze this, we undertake a rather general study of simple groups of rank $6$ in Section 2.3. This culminates with Proposition 3, showing that such groups are so-called $N_{\circ}^{\circ}$-groups in the sense of [DJ16], which we leverage to obtain a contradiction.
@@
We emphasize that the key ingredients of our proof—Propositions 1, 2, and 3—live beyond the specific context of the Theorem and should be interesting in their own right.
Remark.
We wish to highlight that the $L$-hypothesis in the Theorem is only used when we are applying Proposition 1; or more to the point, removing the hypothesis from Proposition 1 would remove it from the Theorem.
Despite some effort trying to free Proposition 1 from the $L$-hypothesis, we were not able to realize it. What is needed is an extension of Corollary 2.6 from actions of $\operatorname{Alt}(n)$ on simple algebraic groups to actions on simple odd-type groups of finite Morley rank. One angle we entertained was to adapt the analysis of actions on degenerate type groups from [AW24]; there, the approach is inductive, driven by the fact that a degenerate-type group acted upon by a Klein $4$-group $K$ is generated by the centralizers of the involutions from $K$ [BBC07, Theorem 5]. But our attempts to generalize this fact to actions on odd-type groups failed to find traction.
There is also the intermediate question of trying to extend Corollary 2.6 to simple $L^{*}$-groups $G$ of odd type. The $L^{*}$-hypotheses assumes all proper definable sections of $G$ are $L$-groups, but makes no claim on $G$ itself. The case of simple $L^{*}$-groups with sufficiently high Prüfer rank has recently been reviewed and clarified by Cherlin: assuming such a $G$ has Prüfer $2$-rank at least $3$ and that there are certain internal restrictions on how $2$-tori act (the NTA2 hypothesis), then either $G$ is algebraic or contains a strongly embedded subgroup [Che24]. As such, a test case for extending Corollary 2.6 to the $L^{*}$-setting might be when $G$ has high Prüfer $2$-rank and a strongly embedded subgroup. Or, instead focusing on low Prüfer $2$-rank, one could try to analyze actions on groups of type $\textbf{CiBo}_{1}$, $\textbf{CiBo}_{2}$, or $\textbf{CiBo}_{3}$ (as defined in Section 2.3) with an initial focus on the minimal simple case.
@@ 2.1. Actions of $\operatorname{Alt}(n)$ on nonsolvable groups
Leveraging the algebraic theory, we solve the conjecture for actions on $L$-groups; this is Proposition 1. Our proof consists to a large degree of simply collecting and organizing known results. As indicated in Corollary 2.5, the bound of $\operatorname{rk}H$ on $n$ is typically rather weak and should be more on the order of $\sqrt{\operatorname{rk}H}$ for generic $n$.
@@
If $n\geq 9$, then $n\leq\operatorname{rk}H$. If the acting group is $\operatorname{Sym}(n)$, then $n\geq 5$ also implies $n\leq\operatorname{rk}H$ unless $n=5$ and $H\cong Z(H)\times\operatorname{PGL}_{2}(F)$ with $\operatorname{char}F=5$.
@@
By [ABC08, Ch. II, Fact 2.25] (generalizing [Hum75, Theorem 27.4]), the acting group is contained in $\operatorname{Inn}(G)\cdot\Gamma$ for $\Gamma$ the group of graph automorphisms relative to a fixed maximal torus and Borel subgroup, so $\operatorname{Alt}(n)$ embeds into $G$ in the cases under consideration. By Corollary 2.5, it only remains to address when $5\leq n\leq 8$ and the acting group is $\operatorname{Sym}(n)$. Let $r=\dim G$. If $r\geq 8$, there is nothing to show, so we may assume $G$ is of type $A(1)$. In particular, $r=3$, and $G=\operatorname{PGL}_{2}(F)$ has no graph automorphisms. Thus, $\operatorname{Sym}(n)$ embeds into $\operatorname{PGL}_{2}(F)$. This is now quite classical. Via the adjoint representation of $\operatorname{PGL}_{2}(F)$, we have a faithful $3$-dimensional representation of $\operatorname{Sym}(n)$, so the only possibility is $n=5=\operatorname{char}F$ (see for example [Dic08]). ∎
@@
Since the Morley rank of an algebraic group $G$ is always greater than or equal to its algebraic dimension, the bound $n\leq\dim G$ appearing in Corollary 2.6 can be replaced by $n\leq\operatorname{rk}G$.
@@
We will denote by $X$ the acting group, $n$ the degree of the permutation group $X$, and $r$ the rank of $H$. Also, set $X^{\prime}=[X,X]=\operatorname{Alt}(n)$, and let $\sigma(H)$ denote the solvable radical of $H$.
We begin by addressing the exceptional case of $n=5$.
Claim 1.
If $X=\operatorname{Sym}(5)$ and $r<5$, then $H\cong Z(H)\times\operatorname{PGL}_{2}(F)$ with $F$ of characteristic $5$.
Proof of Claim.
Since $H$ is nonsolvable and connected groups of Morley rank at most $2$ are solvable [Che79], $r$ is either $3$ or $4$.
First assume $r=3$. Then $H$ is quasisimple with a finite center. Notice that $X$ acts faithfully on $H/Z(H)$; otherwise, the connected group $[X^{\prime},H]$ is contained in $Z(H)$ and hence is trivial (since $Z(H)$ is finite), contradicting faithfulness of the action on $H$. Now, simple groups of rank $3$ are known to be algebraic ([Che79, Fré18]), so Corollary 2.6 applies to $H/Z(H)$. Thus, $H/Z(H)\cong\operatorname{PGL}_{2}(F)$ with $\operatorname{char}F=5$. Moreover, by the theory of central extensions (see [ABC08, Ch. II, Lemma 2.21]), $H$ is algebraic, hence of the form $\operatorname{PGL}_{2}(F)$ or $\operatorname{SL}_{2}(F)$. In either case, $X$ must act by inner automorphisms (see [ABC08, Ch. II, Lemma 2.27]), but $\operatorname{SL}_{2}(F)$ does not embed $\operatorname{Sym}(5)$. Thus, $Z(H)=1$, and $H$ has the desired structure.
Next consider when $r=4$. By [AW24], $H$ contains involutions, and its structure is given by [Wis16]: $H=Z^{\circ}(H)*Q$ with $Q$ quasisimple of rank $3$. Since $Z^{\circ}(H)$ has rank $1$, it is centralized by $X^{\prime}$, so $X$ must act faithfully on $Q$. Our analysis above then shows $Q\cong\operatorname{PGL}_{2}(F)$ in characteristic $5$, which in turn forces $H=Z(H)\times Q$. ∎
For the remainder of the proof, we assume $n\geq 6$ and proceed by contradiction. Thus we assume $r<n$, and we choose $H$ to be a counterexample to Proposition 1 of minimal rank $r$.
@@
Assume $\sigma^{\circ}(H)\neq 1$. Then $\operatorname{rk}(H/\sigma^{\circ}(H))<r$, and since $X$ acts on $H/\sigma^{\circ}(H)$, minimality of $r$ implies that this action is not faithful. Thus $[X^{\prime},H]\leq\sigma^{\circ}(H)$. Now, if we also have $[X^{\prime},\sigma^{\circ}(H)]=1$, then $[X^{\prime},[X^{\prime},H]]=1$, and the three subgroups lemma implies that the group $[X^{\prime},X^{\prime}]=X^{\prime}$ centralizes $H$, against faithfulness of the action on $H$. Thus, $[X^{\prime},\sigma^{\circ}(H)]\neq 1$, so $X$ acts faithfully on $\sigma^{\circ}(H)$.
Then [AW24, Lemma 2.7] implies $\operatorname{rk}(\sigma^{\circ}(H))\geq n-2$ in all cases except possibly when $X=\operatorname{Alt}(9)$. However, if $X=\operatorname{Alt}(9)$ and $\operatorname{rk}(\sigma^{\circ}(H))<7$, we can use that the minimal rank of a faithful connected $\operatorname{Alt}(9)$-module is $7$ (see the main result of [CDW23] and the final remark of Section 3.3, also in [CDW23]) to eliminate this case. Thus $\operatorname{rk}(\sigma^{\circ}(H))\geq n-2$, meaning that $\operatorname{rk}(H/\sigma^{\circ}(H))\leq 2$. Since all connected groups of Morley rank at most $2$ are solvable, this is a contradiction, so $\sigma^{\circ}(H)=1$. ∎
@@
Since $\sigma^{\circ}(H)=1$, we find (as in the proof of Claim 1) that $X$ acts faithfully on $H/\sigma(H)$. Thus $H/\sigma(H)$ is also a counterexample of rank $r$, and the socle of $H$ is as well. Thus we may assume $H$ equals its socle, which is semisimple by [BN94, Theorem 7.8]. ∎
@@
Since $H$ is an $L$-group, each $H_{i}$ is either a simple linear algebraic group over an algebraically closed field or of degenerate type. By the inductive assumption, $n>r\geq k$ (in fact, $n>r\geq 3k$ since the $H_{i}$ are not solvable). As $X^{\prime}$ is simple and permutes the set $\{1,\ldots,k\}$, the only possible action is the trivial one. Thus $X^{\prime}$ normalizes each $H_{i}$. However, $\operatorname{rk}H_{i}\leq r<n$, so $X^{\prime}$ centralizes each $H_{i}$ using Corollary 2.6 in the algebraic case and [AW24] in the degenerate case. This contradicts faithfulness. ∎
@@
If $6\leq n\leq 8$ and $X=\operatorname{Sym}(n)$, we have a contradiction.
@@
By the contradictory assumption, $r\leq 7$. Since connected groups of Morley rank at most $2$ are solvable, $k$ is at most $2$.
We first eliminate the case when $k=2$. If $k=2$ then one of the groups, say $H_{1}$, is of rank $3$ and the other is of rank $4$. As a result $X$ cannot permute these two groups so normalizes each one. Minimality of $r$ implies that the action of $X$ on both components has a nontrivial kernel, which must be $X^{\prime}$. It follows that the action of $X$ on $H$ is not faithful, a contradiction.
Thus, $k=1$ and $H$ is simple. As in Claim 4, Corollary 2.6 and [AW24] provide a contradiction. ∎
The proof of Proposition 1 is now complete. ∎